# Generative AI and Security Operations Center Productivity: Evidence from Live Operations

### v1

@@ 2 Data and Method
Organizations that have consumed Copilot resources are considered treated. To determine treatment times, we use the date the organization first consumed Copilot resources. For each organization we also include the country and industry as organization-level metadata. Finally, we use the number of licenses of the above Microsoft Security products a customer has to conduct our propensity score matching (described below).
@@ 3 Results
Although we do not observe statistically significant differences in the first and second months post adoption, the estimated coefficients trend toward increasing gains over the three-month period. The delay in effect suggests that customers need time to learn about how to best integrate Copilot in their workflows. However, this analysis does not rule out other reasons for the delay such as the time it takes for analysts to adopt Copilot across an organization. The estimates for all three months and their 95% (blue) and 90% (yellow) confidence intervals are given in figure 2
@@ 3.1 Robustness
In general, our results are robust at the 10% significance level for many alternative specifications. As Table 3 shows, the results are largely robust at the 10% significance level with estimates ranging from 22.60% to 31.84%. The only specification that loses statistical significance is when we remove the minimum incident time. However, even there, we easily regain statistical significance by restricting the dataset to organizations that have 20 (compared to 15) observations before and after treatment (effectively removing customers that provide little signal due to lack of observations), with an estimated effect of 30.20% and a p-value of .0625. Furthermore, if we limit the study to large customers (defined by those that have closed at least 100 incidents in the before and after period), we get a similar estimate of the treatment effect (30.78%) but unsurprisingly lose robust statistical significance (p=.265) resulting from the reduction in degrees of freedom. Controlling for organization-specific heterogeneity is a promising direction of future analysis as increased Copilot usage allows for greater statistical power.
@@ 4 Discussion
Moreover, it is even likely that those that are willing to pay for Copilot are the ones that would most benefit from Copilot. In other words, when organizations perform a cost/benefit analysis on the decision to adopt Copilot, only those with the highest benefit would adopt Copilot, since all customers face the same sticker price. This means that, even assuming our estimates are causal, our estimates would overstate the average productivity impact of Copilot for non-adopters.
Considering these limitations, our results provide evidence for the causal impact of Copilot on productivity as measured by MTTR. However, short of performing a randomized control trial on live customer operations or discovering a natural experiment that mimics randomized selection, the associations uncovered in this study are among the most promising signals supporting the positive effect of GAI tools (like Copilot) on the productivity of security operations. Furthermore, the results are statistically significant and robust, so it is clear that those that adopted Copilot observed a productivity gain.
Future research can help control for some of the unobserved confounders. Specifically, moving from a basic difference-in-differences specification to a two-way fixed effects model can help address selection issues by leveraging heterogeneous adoption times and organization fixed effects. The current volume of available data does not provide sufficient power for a two-way fixed effects model. Despite this, the fact that our point estimate when controlling organizational fixed effects shows a similar magnitude of effect as our main result (though not statistically significant) is an encouraging signal for future work.

### v2

@@ 2 Data and Method
Organizations that have consumed Copilot resources are considered treated. To determine treatment times, we use the date the organization first consumed Copilot resources. For each organization we also include the country and industry as organization-level metadata. Finally, we use the number of licenses of the above Microsoft Security products an organization has to conduct our propensity score matching (described below).
@@ 3 Results
Although we do not observe statistically significant differences in the first and second months post adoption, the estimated coefficients trend toward increasing gains over the three-month period. The delay in effect suggests that organizations need time to learn about how to best integrate Copilot in their workflows. However, this analysis does not rule out other reasons for the delay such as the time it takes for analysts to adopt Copilot across an organization. The estimates for all three months and their 95% (blue) and 90% (yellow) confidence intervals are given in figure 2
@@ 3.1 Robustness
In general, our results are robust at the 10% significance level for many alternative specifications. As Table 3 shows, the results are largely robust at the 10% significance level with estimates ranging from 22.60% to 31.84%. The only specification that loses statistical significance is when we remove the minimum incident time. However, even there, we easily regain statistical significance by restricting the dataset to organizations that have 20 (compared to 15) observations before and after treatment (effectively removing organizations that provide little signal due to lack of observations), with an estimated effect of 30.20% and a p-value of .0625. Furthermore, if we limit the study to organizations that have closed at least 100 incidents in the before and after period and perform a two-way fixed effects regression with monthly fixed effects and standard errors clustered [^1] Recent work [cluster] calls into question whether standard errors should be clustered at all. However, we cluster standard errors to be conservative and all of our results remain statistically significant if we do not cluster standard errors. at the organization-time level, we get a similar estimated time savings of 33.69% (p=.039).
@@ 4 Discussion
Moreover, it is even likely that those that are willing to pay for Copilot are the ones that would most benefit from Copilot. In other words, when organizations perform a cost/benefit analysis on the decision to adopt Copilot, only those with the highest benefit would adopt Copilot, since all organizations face the same sticker price. This means that, even assuming our estimates are causal, our estimates would overstate the average productivity impact of Copilot for non-adopters.
Considering these limitations, our results provide evidence for the causal impact of Copilot on productivity as measured by MTTR. However, short of performing a randomized control trial on live organization operations or discovering a natural experiment that mimics randomized selection, the associations uncovered in this study are among the most promising signals supporting the positive effect of GAI tools (like Copilot) on the productivity of security operations. Furthermore, the results are statistically significant and robust, so it is clear that those that adopted Copilot observed a productivity gain.
Future research can help control for some of the unobserved confounders. Specifically, continuing to move from a basic difference-in-differences specification to a two-way fixed effects model can help address selection issues by leveraging heterogeneous adoption times and organization fixed effects. The current volume of available data does not provide sufficient power for a two-way fixed effects model since many organizations only have a small number of incidents. Despite this, the fact that our point estimate when conducting a two-way fixed effects regression on large organizations shows a similar magnitude of effect as our main result is an encouraging signal.